The relative lowering of vapour pressure of an aqueous solution containing non volatile solute is 0.0125. the molality of solution is

Grade:12

2 Answers

Arun
6 years ago
Dear student
 
Relative lowering of vapour pressure = mole fraction of solute (Roult's law)
p- p(s)/p = x2
 
[p - p(s)]/p = w *M/W*m = 0.0125
 
w/m *W = 0.0125/18 = 0.00070
 
Molality = 0.00070 * 1000
= 0.70 M
 
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Relative lowering of vapour pressure = mole fraction of solute (Roult's law)
p- p(s)/p = x2
[p - p(s)]/p = w *M/W*m = 0.0125
w/m *W = 0.0125/18 = 0.00070
Molality = 0.00070 * 1000
= 0.70 M

Thanks and Regards